a survey found that one out of Filipinos says he or she has visited a dentist in any given month. if 12 people are selected at random, find the probability that exactly five will have visited a dentist last month


1
Expert's answer
2021-02-25T01:33:03-0500

The problem is incomplete. Let's use the following assumption: a survey found that one out of five Filipinos says he or she has visited a dentist in any given month.

"p = \frac{1}{5} = 0.2 \\

q = 1 -p = 1 -0.2 = 0.8 \\

n = 12 \\

P(X=x) = \frac{n!}{x!(n-x)!} \times p^x \times (q)^{n-x} \\

P(X=5) = \frac{12!}{5!(12-5)!} \times 0.2^5 \times (0.8)^{12-5} \\

= \frac{8 \times 9 \times 10 \times 11 \times 12}{1 \times 2 \times 3 \times 4 \times 5} \times 0.2^{5} \times 0.8^7 \\

= 0.0531"